# Bounds on the Privacy Amplification of Arbitrary Channels via the Contraction of $f_{\alpha}$-Divergence

### v3

@@ Appendix A Proof of Theorem 8
Finally, we bound the remaining case with the standard Pinsker’s inequality: Since the Rényi-divergence is increasing in $\alpha$ [4], so is $D_{f_{\alpha}}$. Hence the standard Pinsker’s inequlity for the Kullback-Leibler divergence $\text{KL}(P||Q)$ (Rényi-divergence of order $\alpha=1$), $\text{KL}(P||Q)=D_{1}(P||Q)\geq 2\text{TV}(P||Q)^{2}$, will continue to hold for $\alpha>1$. Utilizing (4), this yields
@@ Appendix B Proof of Proposition 3
Since the DPI for Rényi-divergences states that $\eta_{\alpha}(P_{Y|X})\leq 1$ for all $P_{Y|X}$, this then implies $\eta_{\alpha}(P_{Y|X})=1$.
@@
For some very small $\gamma>0$, let $P_{X}^{\gamma}$ be a perturbed single point mass at $x=i$ for some $i\in[N]$, that is,

### v4

@@ Appendix A Proof of Theorem 8
Finally, we bound the remaining case with the standard Pinsker’s inequality: Since the Rényi-divergence is increasing in $\alpha$ [van2014renyi], so is $D_{f_{\alpha}}$. Hence the standard Pinsker’s inequlity for the Kullback-Leibler divergence $\text{KL}(P||Q)$ (Rényi-divergence of order $\alpha=1$), $\text{KL}(P||Q)=D_{1}(P||Q)\geq 2\text{TV}(P||Q)^{2}$, will continue to hold for $\alpha>1$. Utilizing (4), this yields
@@ Appendix B Proof of Proposition 3
Since the DPI for Rényi-divergences states that $\eta_{\alpha}(P_{Y|X})\leq 1$ for all $P_{Y|X}$, this implies $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$.
@@
For some very small $\gamma>0$, let $P_{X}^{\gamma}$ be a uniformly perturbed single point mass at $x=i$ for some $i\in[N]$, that is,